OpenStudy (anonymous):

s=square root 16M simplify then determine the cars speed when skid marks are 121 feet?

OpenStudy (anonymous):

\[\sqrt{16m}\]

OpenStudy (anonymous):

I was able to get \[4\sqrt{m}\]

OpenStudy (cwrw238):

yep thats right now plug in m = 121 to fins the speed

OpenStudy (cwrw238):

*find

OpenStudy (anonymous):

so go like this \[\sqrt{121}\]

OpenStudy (anonymous):

and simplify

OpenStudy (cwrw238):

well 4 * √121

OpenStudy (anonymous):

oh I was forgetting a step

OpenStudy (anonymous):

does 44 look right

OpenStudy (cwrw238):

yes

OpenStudy (cwrw238):

4*11

OpenStudy (anonymous):

yah thanks

OpenStudy (cwrw238):

yw